If 10 gallons of grape fruit juice are added to 40 gallons of mixture, which contains 10% grape juice, then what percent of the resulting mixtue is grape juice?
Answer by lwsshak3(11628)

If 10 gallons of grape fruit juice are added to 40 gallons of mixture, which contains 10% grape juice, then what percent of the resulting mixtue is grape juice?
**
let x=percent of grape juice in resulting mixture
100%*10+10%*40=x%*50
10+4=50x
x=14/50=.28=28%
ans:
percent of grape juice in resulting mixture=28%